Consider

the

following

information

for

a

particular

economy.

Total

population

=

60

million

Number

of

employed

=

30

million

Total

labor

force

=

40

million

Natural

rate

of

unemployment

=

12%

a)

Find

the

total

unemployment

rate

b)

Calculate

the

cyclical

unemployment

rate



1
Expert's answer
2021-10-01T12:23:48-0400

Consider the following information for a particular economy. Total population = 60 million Number of employed = 30 million Total labor force = 40 million Natural rate of unemployment = 12% a) Find the total unemployment rate b) Calculate the cyclical unemployment rate


Answer -

A.

Total labour force = 40 million

Number of employed = 30 million

Number of unemployed = 40 - 30 = 10 million

Unemployment rate = (unemployed / Labour force) × 100

= (10 million / 40 million) × 100

25 %

B.

Cyclical unemployment rate = Unemployment rate - natural rate of unemployment

Unemployment rate = 25%

Natural rate of unemployment = 12%

Cyclical unemployment rate = 25 - 12 = 13 %

(The natural unemployment rate is the sum of frictional unemployment rate and structural unemployment rate.)